Opened 5 years ago

Closed 4 years ago

Last modified 3 years ago

#525 closed flight processing (fixed)

GB13/01, flight day 163a/2014, Bury St Edmunds

Reported by: knpa Owned by:
Priority: alpha 4 high Milestone: 2014 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/GB13_01-2014_163a_Bury_St_Edmunds

Data arrived from ARSF via network transfer on 13/06/2014

GB13/01 Scientific objective: a combination of an ecological model with remotely sensed data can result in an accurate, high resolution map of vegetation communities. Such maps can form the basis to model the distribution of animal species or groups.

GB13/01 Priority: 9

PI: Tim Benton

Notes: Fenix raw files renamed as they were crazy. 1 and 2 are the first two lines on the logsheet. 3 is on 5th line of logsheet and continues onwards normally after that.

Sensors:

Camera (Requested), Archvied 13/05/2015
Fenix (Requested), Archvied 13/05/2015
LiDAR (Requested) , Archvied 13/05/2015
FW (Not Requested), Archvied 13/05/2015

Change History (33)

comment:1 Changed 5 years ago by stgo

Starting navigation processing

comment:2 Changed 5 years ago by stgo

Basestation ppp results:

lat: 52 16 00.50360
long: 0 24 54.96094
ell height: 75.825

Had to correct the elevation mask in ipas tc to 12 degrees to get acceptable solution.

comment:3 Changed 5 years ago by stgo

Beginning hyperspectral processing.

comment:4 Changed 5 years ago by knpa

  • Description modified (diff)

comment:5 Changed 5 years ago by knpa

  • Description modified (diff)

comment:6 Changed 5 years ago by stgo

Beginning RCD processing. Hyperspectral SCT values:
FENIX

-1 1.02
-2 0.98
-3 0.95
-4 0.99
-5 0.94
-6 0.98
-7 0.99
-8 0.97

comment:7 Changed 5 years ago by dap

Beginning LiDAR processing.

comment:8 Changed 5 years ago by tec

Hyperspectral awaiting delivery check.

comment:9 Changed 5 years ago by tec

Beginning RCD

comment:10 Changed 5 years ago by dap

LiDAR roll error value used: -0.005575649 radians.

comment:11 Changed 5 years ago by tec

New Fenix SCT Values

Flightline FENIX
1 1.04
2 1.00
3 0.97
4 1.01
5 0.96
6 1.00
7 1.01
8 0.99
Last edited 5 years ago by tec (previous) (diff)

comment:12 Changed 5 years ago by tec

Fenix waiting delivery check.

comment:13 Changed 5 years ago by tec

RCD waiting for delivery check.

comment:14 Changed 5 years ago by dap

Beginning RCD delivery check.

comment:15 Changed 5 years ago by dap

Finished delivery checking RCD. No problems were found.

comment:16 Changed 5 years ago by dap

LiDAR full waveform delivery creation complete.

comment:17 Changed 5 years ago by benj

Fenix DC started and halted - eight lines in delivery, fourteen on logsheet.

comment:18 Changed 5 years ago by benj

Started Fenix DC again - number of lines is right, auto-generated logsheet is incorrect.

comment:19 Changed 5 years ago by benj

Logsheet has been fixed, fixed APL command in readme to point to view vector file (rather than directory) and removed empty fodis directory. Everything else looks OK, ready to deliver.

comment:20 Changed 5 years ago by dac

Starting delivery check for LiDAR.

comment:21 Changed 5 years ago by dac

LiDAR DC OK, ready to deliver.

Minor comments:

  • First two flight lines large distance from rest.
  • FW data wasn't requested, has been processed anyway - all OK.
  • Extra tiff file in screengrabs for intensity image.
  • No screenshots with vectors.
  • LaTeX syntax errors on LAS1.3 spec hyperlink and quotes around example filename structure (used instead of ) - corrected.
  • compare_dem script produced errors - script / parameterisation problem, DEMs look OK.

comment:22 Changed 5 years ago by dap

  • Deleted .tif screenshot from the screenshots directory.
  • PI was not willing to wait for the vectors screenshot (see forwarded email from tec).
  • The distant flight lines were noticed during processing, but were identified as part of the collected data.
Last edited 5 years ago by dap (previous) (diff)

comment:23 Changed 4 years ago by tec

Archiving
Starting archiving

comment:24 Changed 4 years ago by tec

  • Description modified (diff)
  • Resolution set to fixed
  • Status changed from new to closed

Archiving
Done confirmed as of 13/05/2015 12:10

comment:25 Changed 4 years ago by mark1

Fenix
Fenix data needs reprocessing due to fault with sensor. Incorrect SWIR times recorded in raw files. Have changed tint2 from 14.4 to 2.15 in all raw hdr files.

Fenix data has been reprocessed.

comment:26 Changed 4 years ago by mark1

Fenix

Turned out only 2 lines were affected by the 2.15 SWIR tint (lines 1 + 2), other lines have been set back to 14.4 and reprocessed again.

comment:27 Changed 4 years ago by mark1

Need to make a delivery using the reprocessed lines 3-8 and the original lines 1-2.

comment:28 Changed 4 years ago by dac

Fenix

Copied reprocessed lines to delivery directory - ready for checking.

comment:29 Changed 4 years ago by lah

Fenix re-DC

  • removed empty Eagle & Hawk page from logsheet.
  • lines 3-8 had masked.bil files in level 1 directory. Have removed these. Also removed extra .bil_mask* files
  • underflows look reasonable
  • no zip logfiles (these had been redirected to a temp folder), but zipped files. Remember to remove non-zipped mapped files before delivering.
  • spectra look lovely. Very consistent for each vegetation pixel.

comment:30 Changed 4 years ago by dac

Fenix

Reprocessed data delivered to Tim Benton on USB harddrive 31/07/2015

comment:31 Changed 4 years ago by dac

Fenix Delivery

Hard drive received back

comment:32 Changed 3 years ago by lah

Archiving

Archived 07/10/15. Needs email to Wendy.

Note: See TracTickets for help on using tickets.